The following information pertains to the company:

Work in process, 1/1 (20% complete) 8000 units

Started in January 40000 units

Work in process, 1/31 12000 units

Materials are added at the beginning of the process. If equivalent units of production for conversion using the weighted average method was 39,000, ending work in process at 1/31 must have been

A. 15% complete

B. 30% complete

C. 25% complete

D. 97.5% complete
Answer #1

Correct answer-----------(C) 25% Complete

Working

Statement of Equivalent Units(Weighted average)
Material Conversion cost
Units Complete % Equivalent units Complete % Equivalent units
Transferred                    36,000 100%                    36,000 100%                   36,000
Closing WIP                    12,000 100%                    12,000 25%                     3,000
Total                    48,000 Total                    48,000 Total                   39,000

Units transferred +Units in ending Inventory = equivalent units

Equivalent units = 39000

So 36000+Equivalent Units in ending=39000

Units in Equivalent ending inventory =39000-36000=3000

Whole units in ending inventory = 12000

Equivalent units in ending inventory = 3000

% of completion in ending units = 3000/12000=25%